The volume of a gas-filled balloon is 30.0 L at 40 °C and 153 kPa pressure. What volume will the balloon have at standard temperature and pressure
(273 K and 101.3 kPa)? Round to the nearest tenth. Don't forget the units.
1
Expert's answer
2012-02-24T08:35:24-0500
PV=nRT
n=P1V1/RT1=153 * 30.0/
(273+40)*8.31

n=P2V2/RT2.
V2=nRT2/P2

V2=(153 * 30.0/
(273+40)*8.31)*8.31*273/101.3=39.52 L